OpenStudy (anonymous):

what is the equation of a line with point (-4,0) and a slope of - 3 over 5

OpenStudy (anonymous):

Use point-slope formula for this. \[(y-y_1)=m(x-x_1)\] This gives us y-0=(-3/5)(x+4) From there, you can simplify into slope-intercept if you so choose.

OpenStudy (anonymous):

thanx @stel.la